CAT Chat Minutes

November 15, 2004

 

General Information
Machine Stats – R. Klaffky:  
            Delivered beam:   833 hrs.
            Percentage of scheduled time:   98.06%
            Downtime during period:   16.43 hrs.
            Mean time between faults:   92.53 hrs.

Faults:

# 13   11/09 08:00  To   11/09 13:22     5.38  

  Rad. Mon. Trip      [UES]

  0.75 Spurious Rad Monitor trip at 7-ID, refill, top-up.

# 14   11/09 14:07   To   11/13 02:51    84.74  

 SR-RF4 Trip          [RF]

  0.64 RF-4 power monitor trip, refill, resume top-up.

# 15   11/13 03:30   To   11/15 07:59    52.49  

  Int Dump: End of Period  

  0.00  

Beamline Downtime by APS Groups:   Summary for 11/05-12/2004

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

475

11/09/04

16-ID

Open

2 hr 53 min

ASD-SI

BLP unable to search/secure D-station. Pneumatic Door # 2 open/close LEDs both off and door close button will not activate door closure.

16-ID-D Pneumatic Door # 2 switches required repair.

Beamline Operations restored following APS Management approval.

478

11/12/04

24-ID

Open

18 min

ASD-VAC

BP requested opening the BIV on mezzanine. Confirmed that vacuum pressures were OK and opened BIV valve at 17:34.

 

479

11/10/04

09-ID

Open

46 min

ASD-VAC

BP reported a Serious fault at 1700. BP was trying to open the shutter when fault occurred. Suspected Serious fault #160 (P9CMS1_NOSWITCH) but due to APS network being DOWN (15:57-17:13) - could not confirm what fault occurred. Called Roy Emerson -PSS came to beamline. After network restored - we checked EPICS and saw Serious fault #160 (P9CMS1_NOSWITCH) - Roy reset the fault. Shutter was cycled a number of times without a problem (after fault reset). Fault cleared & shutter opened at 17:46 pm. Roy recommended VACUUM group check out shutter. Trey Leng inspected and will replace seals during Machine Intervention. Submitted Work Request 16308.

 

Safety – B. Glagola:   The first BSL3 experiment ran this weekend; everything went well.

Laser Controlled Area Supervisors Meeting – E. Chang:  The agenda follows for the meeting that will take place Tuesday, November 16 at 1:30 p.m. in 401/B4100.

  • Review of Laser Incident Investigation:   Root causes, Recommendations/Discussion; DOE Video Conference – J. Lang, ASD
  • Restart of Laser Activities at APS:   Issues Identified; J. M. Gibson Participation and Observations – E. Chang, AOD
  • Expert Status Mode – M. Rivers, APS-USR
  • Laser Safety Committee:   Introduction; Distribute draft charter – J. Alicz, XFD

Electrical Safety – Continue to contact the TMS representatives for enrollment in the next safety training being held December 1-3, and 13-14.

TWG Meeting – The upcoming TWG meeting is scheduled to be held at 10:00 a.m. on Thursday, November 18 in 401/A1100.   The topics to be covered are:   Facility Update; HERIX – Concept and Design of a New Spectrometer by H. Sinn, XOR; and The New Ultra-fast Laser Laboratory at Sector 7 by E. Landahl, XOR/MHAT-CAT.